INFRARED INDUCED REACTION OF ATOMIC CHLORINE WITH SOLID PARAHYDROGEN

\underline{PAUL L. RASTON}, DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

The Cl + H_{2} →HCl + H reaction has been studied in solid parahydrogen (pH_{2}) via FTIR spectroscopy of the HCl product. Atomic Cl is generated in solid pH_{2} from the in situ photodissociation of Cl_{2} at 355 nm. The trapped Cl atoms are too cold to react with the surrounding pH_{2} matrix at liquid helium temperatures given the reaction has a barrier of ∼ 5 kcal mol^{−1}. The hydrogen abstraction reaction can not even proceed via quantum mechanical tunneling since the reaction is endothermic by approximately 360 cm^{−1}, which is too great at temperatures of 2-4 K. The reaction can be triggered by allowing IR light from the FTIR spectrometer to impinge upon the Cl atom doped pH_{2} solid. Spectra can be recorded with and without a long bandpass filter present in the IR beam which filters out light above 3900 cm^{−1}. With the filtered light, repeated scans show no evidence for growth of the HCl R_{1}(0) feature and thus no reaction. However, when the filter is removed the HCl absorption grows in intensity signaling the onset of reaction. The integrated absorbance of the HCl feature as a function of time can be fit to a simple first-order rate equation that indicates the Cl atoms have a half-life of 100 minutes under these conditions. We speculate that the reaction is due to solid pH_{2} absorptions that occur above 4100 cm^{−1}, launching H_{2} vibrational excitons (vibrons) into the solid which scatter off the Cl atoms and induce reaction. These experiments are ongoing and the most recent results will be presented at the meeting.